Three uncharged capacitors with equal capacitances are combined in series. The combination is connected to a 5.39-V battery, which charges the capacitors. The charging process involves 4.79 x 104C of charge moving through the battery. Find the capacitance of each capacitor Number
Answer #1

Ans 1/Cnet=1/C+1C+1/C

Cnet=0.33C

Q=Cnet*V

4.76*10^-4=0.33C*5.39

C=2.676*10^-4 F

capacitance of each capacitor=C=2.676*10^-4 F=0.268 mF